### v1

# Answering the Unanswerable Is to Err Knowingly: Analyzing and Mitigating Abstention Failures in Large Reasoning Models

Abstract
Large reasoning models (LRMs) have shown remarkable progress on complex reasoning tasks. However, some questions posed to LRMs are inherently unanswerable, such as math problems lacking sufficient conditions. We find that LRMs continually fail to provide appropriate abstentions when confronted with these unanswerable questions. In this paper, we systematically analyze, investigate, and resolve this issue for trustworthy AI. We first conduct a detailed analysis of the distinct response behaviors of LRMs when facing unanswerable questions. Then, we show that LRMs possess sufficient cognitive capabilities to recognize the flaws in these questions. However, they fail to exhibit appropriate abstention behavior, revealing a misalignment between their internal cognition and external response. Finally, to resolve this issue, we propose a lightweight, two-stage method that combines cognitive monitoring with inference-time intervention. Experimental results demonstrate that our method significantly improves the abstention rate while maintaining the overall reasoning performance.
1 Introduction
Large reasoning models (LRMs), such as GPT-o1 (Jaech et al. 2024) and DeepSeek-R1 (DeepSeek-AI et al. 2025), have demonstrated strong performance on complex reasoning tasks (Xu et al. 2025). By introducing the concept of “thought” and generating longer chains of thought (CoT), LRMs are able to explore diverse reasoning paths while spontaneously reflecting and correcting errors. This enables LRMs to tackle complex tasks with greater depth and flexibility, which is particularly valuable in high-stakes reasoning scenarios.
Figure 1: Examples of abstention failure and success for LRMs on unanswerable questions.
Most current research on LRMs focuses on improving the reasoning process for solving complex problems (Xu et al. 2025), such as reducing redundant reasoning steps to alleviate overthinking (Dang, Huang, and Chen 2025; Yang et al. 2025b) and refining the reasoning process to further enhance reasoning performance (Chen et al. 2025; Liao et al. 2025). However, limited work has considered that not all complex questions are answerable (Kirichenko et al. 2025; Song, Shi, and Zhao 2025). When confronted with noisy and diverse user queries, there will always be cases where LRMs are unable to produce reliable responses, such as the math problems missing sufficient conditions.
As illustrated in Figure 1, when a question is unanswerable due to missing necessary conditions, we prefer LRMs to reason about why it cannot be answered and have the ability to abstain by responding with “I don’t know” (i.e., abstention success), rather than generating a reasoning process with hallucination and arriving at an incorrect answer (i.e., abstention failure). Since model reliability is foundational to user trust, LRMs need to possess both strong reasoning abilities and the capacity to abstain from answering unanswerable questions (Kirichenko et al. 2025; Song, Shi, and Zhao 2025). We first validate the phenomenon that LRMs often struggle to abstain. As shown in Figure 2, we evaluate several LRMs on SUM (Song, Shi, and Zhao 2025), which contains mathematical unanswerable questions. Our results show that most LRMs fail to abstain on more than half of the examples. To address the above issue, we systematically analyze, investigate, and propose solutions to improve the abstention behavior for LRMs.
First, to understand how LRMs fail to abstain, we analyze three types of responses generated by LRMs when faced with unanswerable questions. We find that LRMs exhibit two types of behavior when they fail to abstain. Further, we explore their awareness of unanswerable questions. We conduct analysis on LRMs at both external level (intermediate responses during reasoning) and internal level (latent representations). We find that LRMs possess sufficient cognitive capabilities to recognize flaws in such questions. This reveals a misalignment between internal cognition and external output: a LRM internally realizes that a question is unanswerable, yet still fails to act on this realization and abstain accordingly. Answering the unanswerable is to err knowingly.
Second, we seek to improve the abstention ability of LRMs for unanswerable questions. Our further analysis shows that although LRMs may internally exhibit a tendency to abstain during reasoning, such signals are generally not strong enough to interrupt reasoning and result in abstention. Based on this insight, we propose a method that combines cognitive monitoring with inference-time intervention, aiming to improve LRMs’ ability to abstain from unanswerable questions while preserving their reasoning abilities on answerable ones.
Finally, we conduct extensive experiments with two datasets using LRMs from various model families and scales. Our method enhances LRMs’ ability to abstain from answering unanswerable questions without degrading their reasoning on answerable ones. Furthermore, our experiments reveal that different types of abstention failures benefit from different intervention strategies. Our method achieves significant improvements across all failure types. We will open-source our code to facilitate future research.
Figure 2: Abstention performance comparison of different LRMs for unanswerable questions on the SUM dataset.
2 Related Work
LLM Abstention on Unanswerable Questions.
As the demand for more reliable language models grows, the ability of LLMs to abstain from answering unanswerable questions has become an important evaluation criterion (Yin et al. 2023; Amayuelas et al. 2024; Madhusudhan et al. 2025; Sun et al. 2024; Tomani et al. 2024). LRMs have attracted attention for their strong performance on reasoning tasks (Yang et al. 2025b; Fu et al. 2025). However, their behavior on unanswerable questions has been less studied. Prior work (Kirichenko et al. 2025; Song, Shi, and Zhao 2025) finds that LRMs show weaker abstention ability. In our work, we further provide a more detailed analysis of the different types of outputs produced by LRMs when faced with unanswerable questions. We show that LRMs often have an internal understanding of a problem’s solvability but fail to express it through explicit abstention. Based on this insight, we propose a method to improve their abstention behavior.
Inference-Time Improvements for LRMs.
While LRMs benefit from richer chain-of-thought (CoT) reasoning and achieve strong performance on complex tasks, recent efforts have explored inference-time interventions to further enhance their reasoning accuracy and efficiency (Chen et al. 2025; Dang, Huang, and Chen 2025). For example, (Fu et al. 2025) and (Li et al. 2024) monitor intermediate consistency between reasoning steps to decide when to output the answer. (Yang et al. 2025b) estimates confidence in intermediate steps to decide whether the model has reached a sufficiently certain conclusion. (Liao et al. 2025) uses process-level reward models to guide the model toward better reasoning trajectories. However, the above methods are designed for answerable questions. In our work, we evaluate their effectiveness on unanswerable questions and propose a new inference-time intervention to improve the abstention capability for LRMs.
3 Analysis of Abstention Failure
In this section, we begin by analyzing how LRMs respond to unanswerable math problems. We then investigate LRMs’ awareness of unanswerable questions by examining whether they possess the ability to recognize the unanswerability of such questions, from both internal and external perspectives.
LRMs.
We evaluate five LRMs across diverse model families and scales, including R1-Distill-Llama-8B, R1-Distill-Qwen-7B, R1-Distill-Qwen-14B (DeepSeek-AI et al. 2025), Qwen3-8B, and Qwen3-14B (Yang et al. 2025a).
Datasets.
Following previous work (Song, Shi, and Zhao 2025; Ouyang 2025), we focus on unanswerable math problems, i.e., ill-posed problems that are challenging for LRMs and offer objectively defined criteria for unanswerability. Specifically, we use the Synthetic Unanswerable Math (SUM) (Song, Shi, and Zhao 2025) dataset, which includes diverse problems from AIME (1984–2023), AMC (pre-2023), Omni-MATH (Gao et al. 2025), and Still (Team 2025). The unanswerable problems in SUM are generated based on five criteria: (1) key information deletion, (2) ambiguous key information, (3) unrealistic conditions, (4) unrelated objects, and (5) question deletion. We randomly sample 1,000 problems for our analysis. For each unanswerable case, we use GPT-4o to generate a brief explanation as the ground-truth rationale.
Objective and Prompt.
Given a question, we encourage the LRM to perform normal reasoning when the question is answerable, and to respond with “I don’t know.” along with a corresponding explanation when it is not (i.e., correct abstention). The prompt used in this task is shown in Figure 3.
Figure 3: Prompt used for math problems.
3.1 Reactions to Unanswerable Questions
In this section, we evaluate the reactions of LRMs to unanswerable questions using the prompts in Figure 3 with a maximum token budget of 10,000 per response. We identify three distinct response types, as shown in Figures 4:
Figure 4: Example of different types of response outcomes.
  • Correct abstention: The LRM identifies the question as unanswerable with the given information and explicitly abstains from answering, typically responding with statements such as “I don’t know”.
  • Hallucinated answer: The LRM produces a complete solution by assuming or fabricating missing details not present in the question. As illustrated in Figure 4, the LRM infers a \$9.99 handling charge that is never mentioned in the input in order to compute a final answer.
  • Cognitive fixation: The LRM fails to reach a conclusion within the token limit. It often enters a prolonged reasoning process, stubbornly reformulating or pursuing invalid solution paths without terminating the response, even after recognizing the question is unanswerable.
Figure 5: Distribution of the response types of LRMs on unanswerable math problems.
Figure 5 shows the response type distribution. As model capacity increases, the proportion of correct abstention tends to rise, while those of hallucinated answer and cognitive fixation decrease. A substantial portion of unanswerable questions do not receive correct abstentions. Overall, the results reveal a core limitation: LRMs often fail to abstain from answering unanswerable problems, despite increased model capacity.
3.2 Awareness of Unanswerable Questions
We investigate whether LRMs can recognize unanswerable math problems, probing both their external behavior and internal cognition. Our two-layered analysis assesses (1) behavioral signals: whether LRMs outwardly indicate question answerability, and (2) latent signals: whether answerability is internally encoded during reasoning.
Behavioral Signals of Question Answerability.
Inspired by prior work (Yang et al. 2025b; Chen et al. 2025; Wang et al. 2025) showing that LRMs form and revise intermediate conclusions during reasoning, we insert stopping points into the reasoning trajectory. At each stopping point, we prompt the LRM to (1) directly provide an answer, and (2) to explain why the question is unanswerable (details in the appendix).
Figure 6: The proportion of questions that can respond with a correct abstention in stopping points during the reasoning in the types of “hallucinated answer” and “cognitive fixation”.
Figure 7: The proportion of questions that can provide a correct explanation in stopping points during the reasoning in the types of “hallucinated answer” and “cognitive fixation”.
[…]
Since “I don’t know” is a simple response and may be produced randomly, we additionally prompt the LRMs at each stopping point to explain why the question is unanswerable, to better assess their awareness of unanswerability. We measure the percentage of questions in each category with correct explanations. Figure 7 shows that both failure types yield high percentages of correct explanations.
These findings suggest that even when an LRM fails to abstain, it may still recognize unanswerability during reasoning. Its ability to assess answerability exists but is underused in final decisions.
Figure 8: Classification accuracy of answerable and unanswerable questions at varying stages of the reasoning process.
Latent Signals of Question Answerability.
We conduct a probing-based analysis on the latent representations during reasoning. The linear representation hypothesis posits that high-level concepts, such as language, gender, and truthfulness, are linearly embedded in the latent space of LLMs  (Park, Choe, and Veitch 2024). We hypothesize that question answerability may also be linearly represented during the reasoning process. Inspired by prior work on representation probing (Li et al. 2023; Orgad et al. 2025; Marks and Tegmark 2023; Slobodkin et al. 2023), we train lightweight linear classifiers (probes) on hidden activations from the LRMs’ reasoning trajectory. The goal is to assess whether answerable and unanswerable questions can be distinguished from latent representations during the reasoning process, thereby revealing the presence of answerability-related signals in internal state.
We use the output of the multi-head attention before the residual connection as the input to the probe  (Li et al. 2023):
$x^{c}_{l}=\sum_{h=1}^{H}Q_{l}^{h}\,\mathrm{Att}_{l}^{h}(P_{l}^{h}x_{l}),$ (1)
where $x_{l}$ is the input of layer $l$, $P_{l}^{h}$ projects the input into a head-specific subspace, $Q_{l}^{h}$ maps it back, $\mathrm{Att}$ is the attention operator. The probe is defined as a simple linear classifier: $p_{\theta}(x^{c}_{l})=\sigma(\langle\theta,x^{c}_{l}\rangle)$, where $\theta$ is the trainable weight and $\sigma$ denotes the sigmoid function. One probe is trained per layer.
We sample 2,200 pairs of answerable and unanswerable questions from the SUM dataset (2,000 pairs for training and 200 for validation). For each question, we randomly sample 1,000 token-level activations $x^{c}_{l}$ from the reasoning trajectory, and construct a dataset $\left\{\left(x^{c}_{l},y\right)_{i}\right\}_{i=1}^{N}$, where $y\in\{0,1\}$ indicates question answerability. At inference time, we aggregate the prediction probabilities across all tokens up to the current reasoning step and use the average as the overall answerability prediction. We select the optimal probing layer for each model based on the validation set, and evaluate on the test dataset used for analysis in the previous section.
The results are shown in Figure 8, where we plot the probe’s classification accuracy at various percentages of the reasoning process, using a 0.5 threshold to distinguish unanswerable (prediction probability of probe >> 0.5) from answerable cases. For all LRMs, we observe that the probe’s classification accuracy increases steadily as reasoning progresses, with most accuracies exceeding 0.8 by the end of the trajectory. These results suggest that signals related to question answerability are indeed encoded in the representations of LRMs during reasoning. Although these signals may not always be reflected in the LRMs’ final output behavior, they are implicitly present in the internal computation process.
Figure 9: The confidence of abstention answer: “I don’t know” in different types of response.
Method SUM UMWP
Unanswerable Answerable Unanswerable Answerable
Abstention $\uparrow$ Reason Acc $\uparrow$ Token $\downarrow$ Answer Acc $\uparrow$ Token $\downarrow$ Abstention $\uparrow$ Reason Acc $\uparrow$ Token $\downarrow$ Answer Acc $\uparrow$ Token $\downarrow$
R1-Distill-Llama-8B
Vanilla 16.90% 14.44 5088 61.97 3446 30.67% 26.00 1829 77.67 613
Dynasor-CoT 35.92% 27.82 3084 55.28 2619 39.33% 30.67 958 73.33 495
DEER 24.30% 19.79 4339 60.92 2675 34.11% 28.67 1616 77.67 637
Ours 60.92% 53.17 2419 60.92 3151 54.67% 44.00 1246 77.33 574
R1-Distill-Qwen-7B
Vanilla 21.13% 19.37 4878 69.72 3169 47.67% 43.67 1935 90.30 597
Dynasor-CoT 56.34% 45.89 1869 62.68 2074 64.33% 53.00 763 88.67 486
DEER 28.87% 25.70 3747 63.73 2191 54.33% 49.33 1335 91.67 590
Ours 73.94% 61.86 2247 67.25 3001 77.33% 64.33 1256 90 569
R1-Distill-Qwen-14B
Vanilla 36.97% 33.45 3820 70.42 2671 49.67% 45.00 539 90.00 384
Dynasor-CoT 51.17% 45.18 2622 66.20 2029 50.67% 46.00 504 90.00 384
DEER 39.79% 36.97 3417 63.75 2303 50.33% 46.67 644 90.33 532
Ours 74.30% 62.68 2621 67.96 2541 60.33% 54.33 606 89.67 388
Qwen3-8B
Vanilla 47.18% 41.90 4411 60.92 4245 80.00% 72.33 1906 94.33 1317
Dynasor-CoT 65.61% 58.21 1710 60.27 2190 83.67% 75.00 736 92.00 803
DEER 63.38% 51.17 1902 63.77 1993 80.00% 70.33 479 94.33 474
Ours 75.27% 64.44 2912 61.62 3875 87.33% 79.67 980 93.67 1252
Qwen3-14B
Vanilla 54.22% 48.24 3713 66.55 3768 82.33% 76.67 1279 94.33 877
Dynasor-CoT 66.18% 56.62 1375 63.38 1862 84.00% 76.67 752 92.67 662
DEER 63.90% 56.91 1749 68.18 2192 83.33% 75.67 408 93.00 447
Ours 78.17% 69.01 2311 65.03 3528 92.67% 82.67 959 92.48 848
Table 1: Performance of methods on (un)answerable questions across LRMs. Best scores are marked in bold.
4 Mitigation of Abstention Failure
In this section, we seek to mitigate abstention failures of LRMs when faced with unanswerable questions.
4.1 Motivation
As previously discussed, although LRMs show signs of recognizing unanswerable questions during reasoning, they often fail to abstain in their final answers. We aim to identify the causes of this misalignment and explore ways to fix it.
First, we examine the LRMs’ confidence (Kuhn, Gal, and Farquhar 2023) in abstaining at the stopping point across three output types. We sample an equal number of questions from each output type and compute the average confidence in generating “I don’t know” at the stopping points. Following (Yang et al. 2025b), the confidence score $C$ is computed as
$\begin{aligned} p(a_{t})=P(a_{t}\,|\,H,I,a_{<t}),C=\left(\sum_{i=1}^{n}\max_{a_{t}\in V}p(a_{t})\right)^{\frac{1}{n}},\end{aligned}$ (2)
where $p(a_{t})$ denotes the model’s predicted probability of answer token $a_{t}$, $H$ includes the input prompt and the generated thoughts, $I$ is the prompt used to elicit the answer, $n$ is the number of decoding steps, and $V$ is the vocabulary. As shown in Figure 9, the types of hallucinated answer and cognitive fixation both exhibit lower confidence in abstention compared to the type of correct abstention.
Figure 10: The frequency of abstention responses across stopping points in different types of response outcomes.
Next, we analyze the average frequency of “I don’t know” outputs by the LRMs across all stopping points in the reasoning trajectory. As shown in Figure 10, the hallucinated answer and cognitive fixation types consistently show lower abstention frequencies than correct abstention.
All the findings suggest that while the LRM may recognize a question as unanswerable, it often lacks sufficient confidence to act upon it. The gap between internal awareness and output behavior reveals a key misalignment in LRMs: though they may be aware of unanswerability, their decision process remains biased toward answering rather than abstaining.
4.2 Method
Motivated by the above findings, we propose a method to help LRMs improve their abstention capability. It consists of two key components: cognitive monitoring and inference-time intervention. The goal is to monitor the LRM’s evolving recognition of question unanswerability during reasoning, and to intervene when necessary to guide the LRM toward making an abstention by encouraging abstention behaviors.
Cognitive Monitoring.
The first step of our method aims to identify when the model internally recognizes that a question may be unanswerable. To do so, we track the token-level hidden states generated during inference and segment the reasoning process into semantically coherent units (e.g., clauses or transitions marked by discourse cues such as “wait”). At the end of each segment, we apply a lightweight linear probe, which is trained from our analysis of latent signals of question answerability, to estimate the probability that the current question is unanswerable. If the predicted probability of unanswerability exceeds a threshold, the model transitions to the next stage: inference-time intervention.
Figure 11: The prompt used for inference-time intervention.
Figure 12: Comparison of response type distributions across different methods on unanswerable questions. Proportions are reported across models and datasets, with numbers indicating absolute changes from the vanilla model.
Inference-Time Intervention.
Once the LRM exhibits sufficient internal evidence that a question is unanswerable, we intervene the reasoning process to reinforce this recognition and increase the likelihood of a correct abstention. The intervention is implemented via an instructional guidance prompt (see Figure 11). We append a prompt that restates the possibility that the question may be unanswerable, helping the LRM overcome cognitive fixation and avoid speculative guesses based on missing or unstated information. Inspired by prior work (Yang et al. 2025b; Fu et al. 2025), we also incorporate an early exit strategy to prevent unnecessary continuation of reasoning. It is designed to be minimally intrusive yet semantically salient, encouraging the model to consider abstention as a valid and even preferred option under certain conditions.
5 Experiments
5.1 Setup
Datasets.
We use SUM  (Song, Shi, and Zhao 2025) and UMWP  (Sun et al. 2024). For SUM, we used its test set, which contains 284 unanswerable and 284 answerable questions manually verified. UMWP is derived from four widely used math word problem datasets: SVAMP (Patel, Bhattamishra, and Goyal 2021), MultiArith (Koncel-Kedziorski et al. 2016), GSM8K (Cobbe et al. 2021), and ASDiv (Miao, Liang, and Su 2020). The full dataset consists of 5,200 instances. We sample 600 questions (300 unanswerable and 300 answerable) to form the test set. For the unanswerable questions in UMWP, we also use GPT-4o to generate the reasoning behind their unanswerability as ground truth.
Metrics.
Following previous works (Song, Shi, and Zhao 2025; Kirichenko et al. 2025), we evaluate model performance using the following metrics. (1) Abstention Rate: The proportion of unanswerable questions for which the model correctly abstains from answering by outputting “I don’t know” in accordance with the instruction. (2) Reason Accuracy: The percentage of unanswerable questions for which the model provides the correct rationale for unanswerability. (3) Token Usage: The number of tokens generated by the model during the reasoning process. (4) Answer Accuracy: The proportion of answerable questions for which the model produces the correct final answer.
Baselines and LRMs.
Our method is an inference-time intervention for LRMs without training. There are no directly comparable baselines. We select two baselines based on output consistency (Dynasor-CoT) and confidence (DEER), assuming that correct abstention is the correct answer for unanswerable questions. Dynasor-CoT (Fu et al. 2025) prompts intermediate answers and halts when the same answer appears three times consecutively. DEER (Yang et al. 2025b) monitors sentence-level confidence and exits early once a threshold is met. Besides, the Vanilla method denotes unaltered LRM outputs. For LRMs, we choose five models: R1-Distill-Llama-8B, R1-Distill-Qwen-7B, R1-Distill-Qwen-14B, Qwen3-8B, and Qwen3-14B. Implementation details and full results for all datasets and LRMs are in the appendix.
5.2 Main Results
Table 1 shows the main results. We have the following observations: (1) Our method achieves the highest Abstention Rate and Reason Accuracy on unanswerable questions, demonstrating its strong ability to guide LRMs in recognizing and abstaining from unanswerable inputs. (2) Our method maintains comparable Answer Accuracy on answerable questions. In most settings, accuracy remains close to the vanilla model and sometimes improves slightly, indicating minimal impact on solving answerable tasks. (3) Our method can reduce token usage on unanswerable questions, with an average reduction of 30–50% across different LRMs compared to the vanilla model. It also slightly decreases token consumption on answerable questions, indicating improved reasoning efficiency. (4) We observe a positive correlation between Abstention Rate and Reason Accuracy: as the LRM becomes better at abstaining from unanswerable questions, it also produces more accurate explanations. This suggests that the intervention not only changes the final output but also improves intermediate reasoning quality. (5) Qwen3 models generally outperform distillation-based models in abstention-related metrics. Larger LRMs tend to exhibit stronger abstention capabilities. This trend indicates that both model scale and architecture contribute to reliable unanswerability detection.
5.3 Further Discussions
Impact on Response Type Distribution.
We analyze the changes in the proportions of different response types across methods and datasets, as shown in Figure 12. Our method consistently reduces the hallucinated answers and cognitive fixation outputs. This reduction directly contributes to a substantial increase in the rate of correct abstentions. While both Dynasor-CoT and DEER employ early-exit strategies to mitigate cognitive fixation, we observe that such baselines often lead to a higher proportion of hallucinated answers. Early exits without appropriate guidance may cause the model to make up assumptions or imagined scenarios for giving a definite answer, rather than acknowledging uncertainty. This highlights the importance of combining monitoring with guided intervention to steer LRMs toward proper behavior.
Intervention Effects.
We assess how our intervention influences the LRM’s confidence in abstention and its actual abstention behavior, in order to evaluate whether our method helps bridge the gap between cognitive awareness and abstention behavior. Specifically, at the intervention point identified by our method, we prompt the LRM to generate intermediate outputs both before and after the intervention. We then measure two key indicators: the confidence when producing “I don’t know” responses, and the proportion of questions for which the model outputs “I don’t know”. The results are shown in Table 2. Our method consistently enhances the confidence in generating abstention responses. In addition, the abstention rate also shows corresponding improvements.
Method SUM UMWP
Abst. conf. Abst. rate Abst. conf. Abst. rate
R1-Distill-Llama-8B
Pre-Interv. 79.7 30.1% 84.1 41.5%
Post-Interv. 87.3 ($\uparrow$ 9.4%) 78.1% ($\times$ 2.6) 90.0 ($\uparrow$ 7.0%) 81.4% ($\times$ 1.9)
R1-Distill-Qwen-7B
Pre-Interv. 77.1 24.5% 87.1 50.8%
Post-Interv. 86.8 ($\uparrow$ 12.6%) 80.6% ($\times$ 3.3) 92.1 ($\uparrow$ 5.7%) 71.5% ($\times$ 1.4)
Qwen3-8B
Pre-Interv. 90.9 48.3% 90.6 75.9%
Post-Interv. 98.9 ($\uparrow$ 8.7%) 74.9% ($\times$ 1.5) 98.1 ($\uparrow$ 8.2%) 93.1% ($\times$ 1.2)
Table 2: Results of intervention effects. “Abst. conf.” denotes the average abstention confidence when getting the answer “I don’t know”. “Interv.” is the inference-time intervention.
Further Analysis of Cognitive Monitoring.
We further analyze the cognitive monitoring component of our method and compare our default monitoring strategy based on latent representations with alternative strategies relying on behavioral signals. The behavioral signal approach monitors the LRMs’ intermediate outputs at the end of the paragraph generation phase (e.g., when it reaches a “wait” token), and uses these outputs to determine whether to trigger an intervention. We investigate three variants: The Direct Behavior strategy checks whether the model’s intermediate output is “I don’t know” and triggers an intervention immediately if so. The Consistency strategy triggers intervention only if the model produces “I don’t know” in three consecutive intermediate outputs (inspired by Dynasor-CoT). The Confidence Score strategy triggers intervention when the model outputs “I don’t know” with a confidence score exceeding a predefined threshold (inspired by DEER). As shown in Table 3, all monitoring strategies contribute to the improvement in abstention behavior, showing that cognitive monitoring is generally effective. Among them, the strategy based on latent representation signals achieves the best and most consistent performance across different models and datasets. The Direct Behavior method is simple and works well, but it can be too aggressive and may hurt performance on answerable questions.
Monitoring Signal Unanswerable Ans.
Correct
abstention
$\uparrow$
Hallucinated
answer
$\downarrow$
Cognitive
fixation
$\downarrow$
Acc $\uparrow$
R1-Distill-Llama-8B
Vanilla 16.9 27.8 55.3 61.9
Latent Representation 60.9 15.7 23.4 60.9
Direct Behavior 53.1 20.4 26.5 58.1
Consistency 47.9 23.6 28.5 59.8
Confidence Score 37.0 24.7 38.4 61.3
R1-Distill-Qwen-7B
Vanilla 21.1 25.0 53.9 69.7
Latent Representation 73.9 14.8 11.3 67.3
Direct Behavior 41.6 22.6 35.9 66.8
Consistency 35.7 23.1 41.2 69.3
Confidence Score 31.2 23.9 44.9 69.4
Qwen3-8B
Vanilla 47.2 16.6 36.3 60.9
Latent Representation 75.3 12.4 12.3 61.6
Direct Behavior 67.3 10.9 21.8 60.2
Consistency 61.6 13.0 25.4 61.3
Confidence Score 64.3 10.6 25.2 61.3
Table 3: Comparison of cognitive monitoring strategies.
Ablation Study.
We evaluate two aspects of inference-time intervention: the instructional guidance prompt and the early exit strategy. By removing one component at a time, we analyze how each affects abstention behavior and answer quality. The results are shown in Table 4. For correct abstention, the impact of instructional guidance is greater than that of early exit. The early exit strategy helps reduce the number of cognitive fixation cases. However, without instructional guidance, the proportion of hallucinated answers increases. This again shows that without proper guidance, the model tends to make up conditions and generate unsupported answers. Instructional guidance also has a slight impact on the performance of answerable questions.
Variant Unanswerable Ans.
Correct
abstention
$\uparrow$
Hallucinated
answer
$\downarrow$
Cognitive
fixation
$\downarrow$
Acc $\uparrow$
R1-Distill-Llama-8B
Vanilla 16.9     27.8     55.3   61.9
Ours 60.9     15.7     23.4   60.9
w/o Early Exit 43.3 ($\uparrow$ 26.4)     18.3 ($\downarrow$ 9.5)     38.4 ($\downarrow$ 16.9)   61.9
w/o Instr. Guidance 26.1 ($\uparrow 9.2$)     33.4 ($\uparrow$ 5.6)     40.5 ($\downarrow$ 14.8)   62.3
R1-Distill-Qwen-7B
Vanilla 21.1     25.0     53.9   69.7
Ours 73.9     14.8     11.3   67.3
w/o Early Exit 49.7 ($\uparrow$ 28.6)     16.2 ($\downarrow$ 8.8)     34.2 ($\downarrow$ 19.7)   69.0
w/o Instr. Guidance 43.5 ($\uparrow$ 22.4)     35.2 ($\uparrow$ 10.2)     21.3 ($\downarrow$ 32.6)   70.0
Qwen3-8B
Vanilla 47.2     16.6     36.3   60.9
Ours 75.3     12.4     12.3   61.6
w/o Early Exit 73.6 ($\uparrow$ 26.4)     13.7 ($\downarrow$ 2.9)     12.7 ($\downarrow$ 23.6)   62.7
w/o Instr. Guidance 59.2 ($\uparrow$ 12.0)     30.3 ($\uparrow$ 13.7)     10.6 ($\downarrow$ 25.7)   62.7
Table 4: Ablation results of intervention components across different LRMs on SUM. We report the effect of removing either Instructional Guidance or Early Exit component on three types of responses for unanswerable questions, as well as the accuracy on answerable questions.
6 Conclusion and future work
In this paper, we investigate the failure of LRMs to abstain from answering unanswerable questions, despite having the cognitive ability to detect the unanswerability. We identify a misalignment between the model’s internal cognition and its external response behavior. We propose a lightweight, two-stage method that significantly improves abstention behavior without harming reasoning performance. Future work aims to explore training-time alignment strategies to improve abstention fidelity.
Appendix A Evaluation Details
A.1 Evaluation Details of Three Response Types
In this section, we provide detailed criteria for evaluating the three types of responses introduced in Section 3.1: correct abstention, hallucinated answer, and cognitive fixation. These evaluations are based on model outputs generated in response to an input prompt shown in Figure 3. The criteria are designed for Large Reasoning Models (LRMs).
Correct abstention
A model response is considered a correct abstention if it satisfies the following conditions:
  1. 1.
    The model terminates its reasoning process, indicated by the inclusion of the “</think>” keyword in the output.
  2. 2.
    The model either explicitly includes the phrase “I don’t know” in the answer, or provides an explanation in the format of “Reason {}” indicating that the question is unanswerable.
Hallucinated answer
A model response is labeled as a hallucinated answer if:
  1. 1.
    The model ends its reasoning process by including the “</think>” keyword.
  2. 2.
[…]
Cognitive fixation
A model response is labeled as a cognitive fixation if, under a token budget of 10,000 tokens:
  1. 1.
    The model fails to produce the “</think>” keyword to terminate its reasoning process.
  2. 2.
[…]
A.2 Evaluation Details of Metrics
This section provides detailed descriptions of the evaluation metrics introduced in Section 5 of the main text.
Abstention rate
The proportion of questions that receive the correct abstention among all questions. The criteria for determining a correct abstention are described in Section A.1 of Appendix.
Reason accuracy
The percentage of unanswerable questions for which the model provides the correct rationale for unanswerability. We use the Qwen3-32B model to perform the evaluation, with the judgment prompt shown in Figure 13.
Figure 13: Prompt used for reason accuracy.
Answer accuracy
The proportion of answerable questions for which the model produces the correct final answer. We use the more capable Qwen3-32B model to perform the evaluation, with the specific judgment prompt shown in Figure 14.
Figure 14: Prompt used for answer accuracy.
Figure 15: Comparison of response type distributions on unanswerable questions in the SUM and UMWP datasets. Proportions are reported across methods, with numbers indicating absolute changes from the vanilla model.
Appendix B Implementation Details about Experiments
[…]
Our method consists of two main components. For Cognitive Monitoring, we train a lightweight linear probe on 2,000 pairs of answerable and unanswerable questions sampled from the SUM dataset (Song, Shi, and Zhao 2025), and validate it on an additional 200 pairs. For the reasoning process of each question, we sample 1,000 token-level activations as input to the probe. The linear probe is trained for 75 epochs with a batch size of 16,384 and a learning rate of 3e-5  (Li et al. 2023). We evaluate the classification performance of the probe using the full reasoning process on both the SUM and UMWP test sets. The results are shown in Table 5, where the F1 score is computed with a threshold of 0.5. The results demonstrate that the probe achieves good classification performance and generalizes well across datasets. In our method, the probe is used to predict the answerability of a question based on the generated reasoning content during the inference process. A question is predicted as unanswerable when the probe’s output exceeds a threshold $t$. We set $t$ = 0.6 for the SUM dataset and $t$ = 0.5 for the UMWP dataset.
LRMs
Best
Layer
SUM UMWP
AUROC F1 AUROC F1
R1-Distill-Llama-8B 22 0.879 0.816 0.871 0.809
R1-Distill-Qwen-7B 17 0.912 0.840 0.926 0.867
R1-Distill-Qwen-14B 30 0.908 0.828 0.930 0.879
Qwen3-8B 24 0.925 0.857 0.967 0.919
Qwen3-14B 26 0.936 0.874 0.970 0.918
Table 5: Classification performance of the probe across different LRMs on the SUM and UMWP datasets.
Appendix C Full Results for Further Discussions
Impact on Response Type Distribution.
We analyze the changes in the proportions of different response types for different methods across all LRMs and datasets, as shown in Figure 15. The results are consistent with those presented in Section 5.3 of the main text. Our method consistently reduces the hallucinated answers and cognitive fixation outputs. This reduction directly contributes to a substantial increase in the rate of correct abstentions. While both Dynasor-CoT and DEER employ early-exit strategies to mitigate cognitive fixation, we observe that such baselines often lead to a higher proportion of hallucinated answers. Early exits without appropriate guidance may cause the model to make up assumptions or imagined scenarios for giving a definite answer, rather than acknowledging uncertainty. This highlights the importance of combining monitoring with guided intervention to steer LRMs toward proper behavior.
Intervention Effects.
We assess how our intervention influences the LRM’s confidence in abstention and its actual abstention behavior, in order to evaluate whether our method helps bridge the gap between cognitive awareness and abstention behavior. Specifically, at the intervention point identified by our method, we prompt the LRM to generate intermediate outputs both before and after the intervention. We then measure two key indicators: the confidence when producing “I don’t know” responses, and the proportion of questions for which the model outputs “I don’t know”. The results across all LRMs and datasets are shown in Table 6. The results are consistent with those presented in Section 5.3 of the main text. Our method consistently enhances the confidence in generating abstention responses. In addition, the abstention rate also shows corresponding improvements.
Method SUM UMWP
Abst. conf. Abst. rate Abst. conf. Abst. rate
R1-Distill-Llama-8B
Pre-Interv. 79.7 30.1% 84.1 41.5%
Post-Interv. 87.3 ($\uparrow$ 9.4%) 78.1% ($\times$ 2.6) 90.0 ($\uparrow$ 7.0%) 81.4% ($\times$ 1.9)
R1-Distill-Qwen-7B
Pre-Interv. 77.1 24.5% 87.1 50.8%
Post-Interv. 86.8 ($\uparrow$ 12.6%) 80.6% ($\times$ 3.3) 92.1 ($\uparrow$ 5.7%) 71.5% ($\times$ 1.4)
R1-Distill-Qwen-14B
Pre-Interv. 87.9 47.6% 89.1 65.6%
Post-Interv. 91.9 ($\uparrow$ 4.6%) 77.0% ($\times$ 1.6) 94.9 ($\uparrow$ 6.5%) 84.6% ($\times$ 1.3)
Qwen3-8B
Pre-Interv. 90.9 48.3% 90.6 75.9%
Post-Interv. 98.9 ($\uparrow$ 8.7%) 74.9% ($\times$ 1.5) 98.1 ($\uparrow$ 8.2%) 93.1% ($\times$ 1.2)
Qwen3-14B
Pre-Interv. 91.2 57.8% 94.9 75.6%
Post-Interv. 98.2 ($\uparrow$ 7.7%) 94.7% ($\times$ 1.6) 98.9 ($\uparrow$ 4.1%) 97.3% ($\times$ 1.3)
Table 6: Results of intervention effects. “Abst. conf.” denotes the average abstention confidence when getting the answer “I don’t know”. “Interv.” is the inference-time intervention.
Further Analysis of Cognitive Monitoring.
We analyze the cognitive monitoring component of our method and compare our default monitoring strategy based on latent representations with alternative strategies relying on behavioral signals. The behavioral signal approach monitors the LRMs’ intermediate outputs at the end of the paragraph generation phase (e.g., when it reaches a “wait” token), and uses these outputs to determine whether to trigger an intervention. We investigate three variants: The “direct behavior” strategy checks whether the model’s intermediate output is “I don’t know” and triggers an intervention immediately if so. The “consistency” strategy triggers intervention only if the model produces “I don’t know” in three consecutive intermediate outputs (inspired by Dynasor-CoT). The “confidence score” strategy triggers intervention when the model outputs “I don’t know” with a confidence score exceeding a predefined threshold of 0.95 (inspired by DEER). We conduct experiments on two datasets and five different LRMs, and the results are shown in Table 7. The results are consistent with those presented in Section 5.3 of the main text. All monitoring strategies contribute to the improvement in abstention behavior, showing that cognitive monitoring is generally effective. Among them, the strategy based on latent representation signals achieves the best and most consistent performance across different models and datasets. The direct behavior method is simple and works well, but it can be too aggressive and may hurt performance on answerable questions.
Monitoring Signal SUM UMWP
Unanswerable Ans. Unanswerable Ans.
Correct
abstention
$\uparrow$
Hallucinated
answer
$\downarrow$
Cognitive
fixation
$\downarrow$
Acc $\uparrow$
Correct
abstention
$\uparrow$
Hallucinated
answer
$\downarrow$
Cognitive
fixation
$\downarrow$
Acc $\uparrow$
R1-Distill-Llama-8B
Vanilla 16.9 27.8 55.3 61.9 30.7 53.3 16.0 77.7
Latent Representation 60.9 15.7 23.4 60.9 54.7 41.3 4.0 77.3
Direct Behavior 53.1 20.4 26.5 58.1 51.0 47.0 2.0 75.7
Consistency 47.9 23.6 28.5 59.8 45.3 51.0 3.7 78.3
Confidence Score 37.0 24.7 38.4 61.3 43.3 51.0 5.7 78.3
R1-Distill-Qwen-7B
Vanilla 21.1 25.0 53.9 69.7 47.7 36 16.3 90.3
Latent Representation 73.9 14.8 11.3 67.3 77.3 20.3 2.3 90.0
Direct Behavior 41.6 22.6 35.9 66.8 61.3 32.0 6.7 90.0
Consistency 35.7 23.1 41.2 69.3 57.3 34.7 8.0 91.0
Confidence Score 31.2 23.9 44.9 69.4 54.3 34.3 11.3 90.3
R1-Distill-Qwen-14B
Vanilla 36.9 25.0 38.0 70.4 49.7 49.7 0.7 90.0
Latent Representation 74.3 15.9 9.9 67.9 60.3 39.7 0.0 89.7
Direct Behavior 67.3 19.0 13.7 64.4 53.7 46.0 0.3 90.0
Consistency 62.3 21.1 16.6 65.9 50.0 49.7 0.3 90.0
Confidence Score 53.9 22.2 23.9 69.4 52.0 47.7 0.3 91.0
Qwen3-8B
Vanilla 47.2 16.6 36.3 60.9 80.0 18.7 1.3 94.3
Latent Representation 75.3 12.4 12.3 61.6 87.3 12.3 0.3 93.7
Direct Behavior 67.3 10.9 21.8 60.2 91.3 8.3 0.3 90.0
Consistency 61.6 13.0 25.4 61.3 87.3 12.3 0.3 92.7
Confidence Score 64.3 10.6 25.2 61.3 90.3 9.3 0.3 91.3
Qwen3-14B
Vanilla 54.2 17.9 27.8 66.6 82.3 17.3 0.3 94.3
Latent Representation 78.2 9.5 12.3 65.0 92.7 7.3 0.0 92.5
Direct Behavior 74.3 9.5 16.2 64.0 92.0 8.0 0.0 90.7
Consistency 67.9 12.3 19.7 66.2 87.3 12.7 0.0 93.0
Confidence Score 71.5 9.9 18.7 65.1 92.0 8.0 0.0 91.6
Table 7: Comparison of different cognitive monitoring strategies across different LRMs on SUM and UMWP datasets.
Ablation Study.
We evaluate two aspects of inference-time intervention: the instructional guidance prompt and the early exit strategy. By removing one component at a time, we analyze how each affects abstention behavior and answer quality. The results, shown in Table 8, are consistent with those presented in Section 5.3 of the main text. For correct abstention, the impact of instructional guidance is greater than that of early exit. The early exit strategy helps reduce the number of cognitive fixation cases. However, without instructional guidance, the proportion of hallucinated answers increases. This again shows that without proper guidance, the model tends to make up conditions and generate unsupported answers. Instructional guidance also has a slight impact on the performance of answerable questions.
Variant SUM UMWP
Unanswerable Ans. Unanswerable Ans.
Correct
abstention
$\uparrow$
Hallucinated
answer
$\downarrow$
Cognitive
fixation
$\downarrow$
Acc $\uparrow$
Correct
abstention
$\uparrow$
Hallucinated
answer
$\downarrow$
Cognitive
fixation
$\downarrow$
Acc $\uparrow$
R1-Distill-Llama-8B
Vanilla 16.9     27.8     55.3   61.9 30.7 53.3 16.0 77.7
Ours 60.9     15.7     23.4   60.9 54.7 41.3 4.0 77.3
w/o Early Exit 43.3 ($\uparrow$ 26.4)     18.3 ($\downarrow$ 9.5)     38.4 ($\downarrow$ 16.9)   61.9 45.7 ($\uparrow$ 15.0) 48.7 ($\downarrow$ 4.6) 5.7 ($\downarrow$ 10.3) 77.0
w/o Instr. Guidance 26.1 ($\uparrow 9.2$)     33.4 ($\uparrow$ 5.6)     40.5 ($\downarrow$ 14.8)   62.3 41.0 ($\uparrow$ 10.3) 55.3 ($\uparrow$ 2.0) 3.7 ($\downarrow$ 12.3) 78.0
R1-Distill-Qwen-7B
Vanilla 21.1     25.0     53.9   69.7 47.7 36.0 16.3 90.3
Ours 73.9     14.8     11.3   67.3 77.3 20.3 2.3 90.0
w/o Early Exit 49.7 ($\uparrow$ 28.6)     16.2 ($\downarrow$ 8.8)     34.2 ($\downarrow$ 19.7)   69.0 67.0 ($\uparrow$ 19.3) 26.0 ($\downarrow$ 10.0) 7.0 ($\downarrow$ 9.3) 90.3
w/o Instr. Guidance 43.5 ($\uparrow$ 22.4)     35.2 ($\uparrow$ 10.2)     21.3 ($\downarrow$ 32.6)   70.0 58.0 ($\uparrow$ 10.3) 40.7 ($\uparrow$ 4.7) 1.3 ($\downarrow$ 15.0) 91.0
R1-Distill-Qwen-14B
Vanilla 36.9     25.0     38.0   70.4 49.7 49.7 0.7 90.0
Ours 74.3     15.9     9.9   67.9 60.3 39.7 0.0 89.7
w/o Early Exit 63.4 ($\uparrow$ 26.5)     20.1 ($\downarrow$ 4.9)     16.6 ($\downarrow$ 21.4)   69.4 55.0 ($\uparrow$ 5.3) 44.7 ($\downarrow$ 5.0) 0.3 ($\downarrow$ 0.4) 90.0
w/o Instr. Guidance 46.8 ($\uparrow$ 9.9)     35.2 ($\uparrow$ 10.2)     17.9 ($\downarrow$ 20.1)   70.7 51.0 ($\uparrow$ 1.3) 49.0 ($\downarrow$ 0.7) 0.0 ($\downarrow$ 0.7) 90.0
Qwen3-8B
Vanilla 47.2     16.6     36.3   60.9 80.0 18.7 1.3 94.3
Ours 75.3     12.4     12.3   61.6 87.3 12.3 0.3 93.7
w/o Early Exit 73.6 ($\uparrow$ 26.4)     13.7 ($\downarrow$ 2.9)     12.7 ($\downarrow$ 23.6)   62.7 86.3 ($\uparrow$ 6.3) 13.3 ($\downarrow$ 5.4) 0.3 ($\downarrow$ 1.0) 94.0
w/o Instr. Guidance 59.2 ($\uparrow$ 12.0)     30.3 ($\uparrow$ 13.7)     10.6 ($\downarrow$ 25.7)   62.7 80.5 ($\uparrow$ 0.5) 19.2 ($\uparrow$ 0.5) 0.3 ($\downarrow$ 1.0) 94.3
Qwen3-14B
Vanilla 54.2     17.9     27.8   66.6 82.3 17.3 0.3 94.3
Ours 78.2     9.5     12.3   65.0 92.7 7.3 0.0 92.5
w/o Early Exit 73.6 ($\uparrow$ 19.4)     11.9 ($\downarrow$ 6.0)     14.4 ($\downarrow$ 13.4)   65.0 92.0 ($\uparrow$ 9.7) 8.0 ($\downarrow$ 9.3) 0.0 ($\downarrow$ 0.3) 92.7
w/o Instr. Guidance 63.0 ($\uparrow$ 8.8)     23.9 ($\uparrow$ 6.0)     13.0 ($\downarrow$ 14.8)   66.9 85.0 ($\uparrow$ 2.7) 15.0 ($\downarrow$ 2.3) 0.0 ($\downarrow$ 0.3) 94.0
Table 8: Ablation results of intervention components across different LRMs on SUM and UMWP datasets. We report the effect of removing either Instructional Guidance or Early Exit component on three types of responses for unanswerable questions, as well as the accuracy on answerable questions.